OpenStudy (anonymous):

Find average value f(theta)= sec^2(theata/2), [0/pi]

OpenStudy (anonymous):

\[sec^2(\theta/2)\]

OpenStudy (anonymous):

\[\frac{1}{\pi -0}\int_0^{\pi } \text{Sec}^2\left(\frac{x}{2}\right) \, dx\]

OpenStudy (anonymous):

you need help integrating that?

OpenStudy (anonymous):

would it be\[1/\Pi\] ?

OpenStudy (anonymous):

The integral doesn't converge http://www.wolframalpha.com/input/?i=Integrate%5BSec^2%5Bx%2F2%5D%2C {x%2C0%2CPi}%5D

OpenStudy (anonymous):

http://tinyurl.com/3qhwlkk